Synthetic division: we can use a shortcut to long division of polynomials called synthetic division if the divisor d(x) is in the form d(x)=x-c.
a. if the divisor is x-8 then c=_____
b. if the divisor is x+6 then c =_______
i need help with this explanation before i can continue my math homework in this subject, thank you

Answer by Boreal(15235)

x-c=x-8; c=8
x-c=x+6; c=-6
x^2 +6x +8 has roots (x+4) and (x+2) both equaling 0. x=-4, x=-2
-2; 1;;;6;;;;8
;;;1;;4;;;-8
;;;;;(x+4)
you take the opposite value of the factor, what you get when you set it equal to zero, the true root, and use it as a divisor.
So for x-c, you use c.
x-c=0; x=c